A solution is made by diluting 25mL of 0.34 mol/L calcium nitrate, Ca(NO3)2 (aq) solution to 100 Ml. Calculate the following concentrations for the solution:
a. the concentration of calcium nitrate
b. the concentration of nitrate ions
1
Expert's answer
2013-01-08T11:59:42-0500
The amount of calcium nitrate = 25*0.34/1000 = 0.0085 mol
New volume of solution is 25 + 100 = 125 ml
a) New concentration of calcium nitrate = 0.0085 mol / 125 ml or 0.068 mol/L (0.068M)
b)The concentration of nitrate ions is twice more than concentration of calcium nitrate (Сa(NO3)2-----2NO3-) = 0.136 M